# Multiscale Quantile Regression with Local Error Control

### v2

@@ A.1 Technical results
$\tilde{\theta}=\begin{cases}\min\{Z_{i}\,|\,Z_{i}>0,\,i=1,\dots,n\}/2,&\quad\text{if }\exists\,Z_{i}>0,\\ 1,&\quad\text{otherwise.}\end{cases}$
Then $\tilde{\theta}>0$. Moreover, by the definition of $\tilde{\theta}$, for any $l\in\mathbb{N}$ the events $\{Z_{i}\leq 0\}$ and $\{Z_{i}\leq\tilde{\theta}/l\}$ are equivalent. Define the stopping time
$T_{1}=\max\left\{j\,|\,T_{[0/n,j/n)}(W,0)\leq q_{\alpha}(j)\right\}.$
Set $I_{0}=[0/n,T_{1}/n)$. Then by definition of $T_{1}$ we have $T_{\accentset{\circ}{I_{0}}}(W,0)\leq q_{\alpha}(T_{1})$. We defined the modification $\widetilde{W}_{i}=\mathbbm{1}_{\{Z_{i}\leq\tilde{\theta}\}}$ for all $i/n\in I_{0}$, then $W_{i}=\widetilde{W}_{i}$. Thus,
$T_{\accentset{\circ}{I_{0}}}(\widetilde{W},\tilde{\theta})=T_{\accentset{\circ}{I_{0}}}(W,0)\leq q_{\alpha}(T_{1}),$
@@
which indicates that $\tilde{\theta}$ fulfils the multiscale side-constrain on $\accentset{\circ}{I_{0}}$ and therefore, we can estimate ${f}|_{I_{0}}$ with $\tilde{\theta}$. We continue this procedure until obtain an estimator for whole ${f}$. More precisely, for any $k\in\mathbb{N}$, let
$T_{k}=\max\left\{j\,|\,T_{[T_{k-1}/n,j/n)}(W,0)\leq q_{\alpha}(j-T_{k-1})\right\},$
and $I_{k-1}=[T_{k-1}/n,T_{k}/n)$. The restriction ${f}|_{I_{k-1}}$ can be estimated by $\tilde{\theta}/k$. Finally, let
$\widetilde{K}=\max\{k\,|\,T_{k}<n\},$
and estimation is given by
$\widetilde{f}=\sum_{k=0}^{\tilde{K}}\frac{\tilde{\theta}}{k+1}\mathbbm{1}_{I_{k}}.$
@@
Since $\tilde{f}$ is piecewise different, it has exactly $\widetilde{K}$ change points. Thus, applying Proposition 7 we have
$\mathbb{P}\left(\widehat{K}\geq k\right)\leq\mathbb{P}\left(\widetilde{K}\geq k\right)\leq\alpha^{k},$
and
$\mathbb{E}\left(\widehat{K}\right)\leq\mathbb{E}\left(\widetilde{K}\right)\leq\frac{\alpha}{1-\alpha}.$
A.2 Proof of Theorems 14
@@
$2(p-q)^{2}\leq p\log\left(\frac{p}{q}\right)+(1-p)\log\left(\frac{1-p}{1-q}\right)\leq\frac{(p-q)^{2}}{q(1-q)}.$
@@
The first inequality follows from Pinsker’s inequality (Tsybakov, , 2009, Lemma 2.5) and (Vanegas et al., , 2022, Theorem A.3). For the second one, we apply the basic inequality $x(x-1)\geq x\log(x)$ for $x>0$, then
$\frac{p}{q}\log\left(\frac{p}{q}\right)$ $\leq\frac{p}{q}\left(\frac{p}{q}-1\right)$
$\frac{1-p}{1-q}\log\left(\frac{1-p}{1-q}\right)$ $\leq\frac{1-p}{1-q}\left(\frac{1-p}{1-q}-1\right)$
@@
$T_{I}(Z,\theta)\;=\;\underset{J\subseteq{I}}{\max}\sqrt{2L_{J}(Z,\theta)}-\sqrt{2\log\left(\frac{e\lvert I\rvert}{\lvert J\rvert}\right)},$
where
$L_{J}(Z,\theta)=\lvert J\rvert\left[\overline{W}_{J}\log\left(\frac{\overline{W}_{J}}{\beta}\right)+(1-\overline{W}_{J})\log\left(\frac{1-\overline{W}_{J}}{1-\beta}\right)\right],$
$\overline{W}_{J}=\left(\sum_{i\in J}W_{i}\right)/\lvert J\rvert$ and $W_{i}=\mathbbm{1}_{\{Z_{i}\leq\theta\}}\overset{i.i.d.}{\sim}{\rm Ber}(\beta)$ under $H_{I}$. Since $T_{I}(Z,\theta)$ is determined by $W$, in this proof we use $T_{I}(W)$ instead. For any $J\subseteq I$, we define $a_{J}\in\mathbb{R}^{\lvert I\rvert}$ by
$a_{J,k}=\begin{cases}\frac{1}{\sqrt{\lvert J\rvert}},&\quad\text{if }k\in J,\\ 0,&\quad\text{otherwise},\end{cases}$
where $a_{J,k}$ denotes the $k$-th component of $a_{J}$. Set $A=\{a_{J}\,|J\subseteq I\}$ and $\lambda_{a}=\sqrt{2\log en\lVert a\rVert_{\infty}^{2}}$. We further define $R_{I}(W)$ and $S_{I}(W)$ by
$R_{I}(W)$ $=\underset{a\in A}{\max}\sqrt{2}\lvert a^{\top}(W-\beta)\rvert-\lambda_{a}$
$S_{I}(W)$ $=\underset{a\in A}{\max}\sqrt{\frac{2}{\beta(1-\beta)}}\lvert a^{\top}(W-\beta)\rvert-\lambda_{a}.$
Then by Lemma 9 we have
$R_{I}\leq T_{I}\leq S_{I}.$
For any $x,y\in\{0,1\}^{\lvert I\rvert}$ and weight $a\in\mathbbm{R}_{\geq 0}^{\lvert I\rvert}$, the weighted Hamming distance $d_{a}$ is given by
$d_{a}(x,y)=\sum_{i=1}^{\lvert I\rvert}a_{i}\mathbbm{1}_{\{x_{i}\neq y_{i}\}}.$
We show that $T_{I}$ has the following Lipschitz property for the distance $d_{a}$ with Lipschitz constant $\sqrt{2}$: For any $x\in\{0,1\}^{\lvert I\rvert}$, there exists a $a=a(x)\in\mathbbm{R}_{\geq 0}^{\lvert I\rvert}$ with Euclidean norm $\lVert a\rVert=1$ such that
$T_{I}(x)-T_{I}(y)\leq Ld_{a}(x,y),\quad\forall y\in\{0,1\}^{\lvert I\rvert}.$
For any $x,y\in\{0,1\}^{\lvert I\rvert}$, let $\tilde{a}=\tilde{a}(x)$ denote the maximizer of $S_{I}(x)$
$\,T_{I}(x)-T_{I}(y)$
$\leq$ $\,S_{I}(x)-R_{I}(y)$
@@
where the second last equality follows the fact $x,y\in\{0,1\}^{\lvert I\rvert}$. Let $M_{I}$ denote the median of $T_{I}$. By Frick et al., (2014); Dümbgen and Spokoiny, (2001); Dümbgen et al., (2006) the random sequence $T_{I}$ converges in distribution a almost surely finite random variable $M$ as $\lvert I\rvert\to\infty$, thus median of $M$ is finite and $C\coloneqq\sup_{\lvert I\rvert\in\mathbb{N}}M_{I}<\infty$. Applying (Giné and Nickl, , 2016, Corollary 3.2.5) we have for any $t>0$
$\mathbb{P}(T_{I}-M_{I}\geq t)\leq 2\exp{\left(-t^{2}/8\right)}.$
By choosing $t=2\sqrt{2\log(2/\alpha)}$ we have for any $\lvert I\rvert\in\mathbb{N}$
$q_{\alpha}(\lvert I\rvert)\leq M_{I}+2\sqrt{2\log\left(\frac{2}{\alpha}\right)}\leq C+2\sqrt{2\log\left(\frac{2}{\alpha}\right)}.$
@@
$I_{k}=\left(\tau_{k}-\frac{\lambda_{k}}{2},\tau_{k}+\frac{\lambda_{k}}{2}\right),$
@@
$I_{k}^{+}=\{t\in I_{k}\,|\,f(t)=\theta_{k}^{+}\},\text{ and }I_{k}^{-}=\{t\in I_{k}\,|\,f(t)=\theta_{k}^{-}\}.$
Assume that $\widehat{K}<K$ and let $\widehat{I}_{0},\dots,\widehat{I}_{\widehat{K}}$ denote the estimated segments. Since $\widehat{K}<K$, there exists a interval $I_{k}$ such that $I_{k}$ is completely contained in some estimated segment, say $\widehat{I}_{*}$. This can be verified by contradiction: Suppose that there is no such interval, then any interval $I_{k}$ contains at least one estimated change point. Since $I_{k}$’s are disjoint, then we have $\widehat{K}\geq K$, which contradicts to the underestimation assumption. Thus, there exists a $I_{k}$ is contained in some estimated segment $\widehat{I}_{*}$, i.e., no change point is detected on $I_{k}$.
@@
$C_{k}=\Bigg{\{}\exists c\in\mathbb{R}\Bigg{|}\sqrt{2L_{I_{k}^{+}}(Z,c)}-\sqrt{2\log\left(\frac{e\lvert\widehat{I}_{*}\rvert}{\lvert I_{k}^{+}\rvert}\right)}\leq q_{\alpha}(\lvert\widehat{I}_{*}\rvert)\text{ and }\sqrt{2L_{I_{k}^{-}}(Z,c)}-\sqrt{2\log\left(\frac{e\lvert\widehat{I}_{*}\rvert}{\lvert I_{k}^{-}\rvert}\right)}$ $\leq q_{\alpha}(\lvert\widehat{I}_{*}\rvert)\Bigg{\}}$
@@
$C_{k}^{+}$ =$=\left\{\exists c\geq\theta_{k}^{+}-\frac{\delta_{k}}{2}\bigg{|}\sqrt{2L_{I_{k}^{+}}(Z,c)}-\sqrt{2\log\left(\frac{e\lvert\widehat{I}_{*}\rvert}{\lvert I_{k}^{+}\rvert}\right)}\leq q_{\alpha}(\lvert\widehat{I}_{*}\rvert)\right\},$
$C_{k}^{-}$ =$=\left\{\exists c\leq\theta_{k}^{-}+\frac{\delta_{k}}{2}\bigg{|}\sqrt{2L_{I_{k}^{-}}(Z,c)}-\sqrt{2\log\left(\frac{e\lvert\widehat{I}_{*}\rvert}{\lvert I_{k}^{-}\rvert}\right)}\leq q_{\alpha}(\lvert\widehat{I}_{*}\rvert)\right\}.$
Obviously, we have $C_{k}\subseteq C_{k}^{+}\cup C_{k}^{-}$. Note that $C_{k}^{+}$ and $C_{k}^{-}$ are independent. We obtain
$\mathbb{P}(C_{k})\leq 1-(1-\mathbb{P}(C_{k}^{+}))(1-\mathbb{P}(C_{k}^{-})).$
We provide an upper bound of $\mathbb{P}(C_{k}^{-})$, then the probability $\mathbb{P}(C_{k}^{+})$ can be upper bounded in a similar way. For any $i\in I_{k}^{-}$, we have $F_{i}^{-1}(\beta)=\theta_{k}^{-}$. Further set $F^{-}_{\min,k}=\min_{i\in I_{k}^{-}}F_{i}$ and $\bar{F}_{k}^{-}=\left(\sum_{i\in I_{k}^{-}}F_{i}\right)/\lvert I_{k}^{-}\rvert$. Let
$\zeta_{I_{k}^{-}}=\inf\left\{x\in\mathbb{R}\Bigg{|}\sum_{i\in I_{k}^{-}}\mathbbm{1}_{\{Z_{i}\leq x\}}\geq\lvert I_{s}^{-}\rvert\beta\right\}$
be the empirical quantile of the observations on $I_{k}^{-}$. For all $c\geq\theta_{k}^{-}+\delta_{k}/2$, if $\zeta_{I_{k}^{-}}\leq\theta_{k}^{-}+\delta_{k}/2$, then
@@
where $\overline{W}(Z,x)=\sum_{i\in I_{k}^{-}}\mathbbm{1}_{\{Z_{i}\leq x\}}/\lvert I_{k}^{-}\rvert$. Note that the function
$g:x\mapsto\lvert I_{k}^{-}\rvert\left(x\log\left(\frac{x}{\beta}\right)+(1-x)\log\left(\frac{1-x}{1-\beta}\right)\right)$
is strictly convex with minimum $g(\beta)=0$, hence $g$ is strictly increasing on $(\beta,1)$. Thus, by (22)
$L_{I_{k}^{-}}(Z,c)=g\left(\overline{W}(Z,c)\right)\geq g\left(\overline{W}\left(Z,\theta_{k}^{-}+\frac{\delta_{k}}{2}\right)\right)=L_{I_{k}^{-}}\left(Z,\theta_{k}^{-}+\frac{\delta_{k}}{2}\right).$
Therefore,
$\mathbb{P}(C_{k}^{-})\leq$ $\,\mathbb{P}\left(C_{k}^{-}\cap\left\{\zeta_{I_{k}^{-}}\leq\theta_{k}^{-}+\frac{\delta_{k}}{2}\right\}\right)+\mathbb{P}\left(\zeta_{I_{k}^{-}}>\theta_{k}^{-}+\frac{\delta_{k}}{2}\right)$
$\leq$ $\,\mathbb{P}\left(\sqrt{2L_{I_{k}^{-}}\left(Z,\theta_{k}^{-}+\frac{\delta_{k}}{2}\right)}-\sqrt{2\log\left(\frac{e\lvert\widehat{I}_{*}\rvert}{\lvert I_{k}^{-}\rvert}\right)}\leq q_{\alpha}(\lvert\widehat{I}_{*}\rvert)\right)+\mathbb{P}\left(\zeta_{I_{k}^{-}}>\theta_{k}^{-}+\frac{\delta_{k}}{2}\right)$
$\leq$ $\,\mathbb{P}\left(L_{I_{k}^{-}}\left(Z,\theta_{k}^{-}+\frac{\delta_{k}}{2}\right)\leq\frac{\left(q_{\alpha}(\lvert\widehat{I}_{*}\rvert)+\sqrt{2\log\left(\frac{e\lvert\widehat{I}_{*}\rvert}{\lvert I_{k}^{-}\rvert}\right)}\right)^{2}}{2}\right)+\mathbb{P}\left(\zeta_{I_{k}^{-}}>\theta_{k}^{-}+\frac{\delta_{k}}{2}\right)$
$\leq$  $\,2\exp\left(-\frac{\left(\sqrt{2n\lambda_{k}}\xi_{\bar{F}^{-}_{\beta}}(\delta_{k}/2)-q_{\alpha}(\lvert\widehat{I}_{*}\rvert)-\sqrt{2\log\left(\frac{e\lvert\widehat{I}_{*}\rvert}{\lvert I_{k}^{-}\rvert}\right)}\right)^{2}_{+}}{2}\right)+2\exp\left(-n\lambda_{k}\xi_{F^{-}_{\min,\beta}}(\delta_{k}/2)^{2}\right),$
where the last two inequalities follow from (Vanegas et al., , 2022, Theorem A.2 and A.3), respectively. Note that the facts $\xi_{k}\leq\xi_{F^{-}_{\min,k},\beta}(\delta_{k}/2)\leq\xi_{\bar{F}_{k}^{-},{\beta}}(\delta_{k}/2)$, $q_{\alpha}(\lvert\widehat{I}_{*}\rvert)\leq\max_{m\leq n}q_{\alpha}(m)$ and $\lvert\widehat{I}_{*}\rvert\leq n$ we have
$\mathbb{P}(C_{k})\leq 1-(1-\mathbb{P}(C_{k}^{+}))(1-\mathbb{P}(C_{k}^{-}))\leq 1-\gamma_{n,k}^{2},$
where $\gamma_{n,k}$ is given by
$\gamma_{n,k}=1-2\exp{\left(-n\lambda_{k}\xi_{k}^{2}\right)}-2\exp{\left(-\left(\sqrt{n\lambda_{k}}\xi_{k}-\frac{1}{\sqrt{2}}\underset{m\leq n}{\max}q_{\alpha}(m)-\sqrt{\log\frac{2e}{\lambda_{k}}}\right)_{+}^{2}\right)}.$
Note that if all of these $K$ change-points are detected, then $\widehat{K}\geq K$, i.e.,
@@
which proves the first claim. By the definitions of $\Lambda$ and $\Omega$ in (8) we have for all $k=1,\dots,K$,
$\gamma_{n,k}^{2}\geq\Bigg{(}1-2\exp\left(-\left(\sqrt{n}\Omega-\frac{1}{\sqrt{2}}\underset{m\leq n}{\max}q_{\alpha}(m)-\sqrt{\log\left(\frac{2e}{\Lambda}\right)}\right)_{+}^{2}\right)-2\exp\left(-n\Omega^{2}\right)\bigg{)}^{2}\eqqcolon\gamma_{n}.$
Thus, applying Bernoulli’s inequality we have
$\mathbb{P}(\widehat{K}<K)\leq 1-\gamma_{n}^{K}\leq 4K\Bigg{(}\exp\left(-\left(\sqrt{n}\Omega-\frac{1}{\sqrt{2}}\underset{m\leq n}{\max}q_{\alpha}(m)-\sqrt{\log\left(\frac{2e}{\Lambda}\right)}\right)_{+}^{2}\right)+\exp\left(-n\Omega^{2}\right)\Bigg{)}.$

### v3

@@ A.1 Technical results
$\tilde{\theta}=\begin{cases}\min\{Z_{i}:Z_{i}>0,\,i=1,\dots,n\}/2,&\qquad\text{if there is a }\,Z_{i}>0,\\ 1,&\qquad\text{otherwise.}\end{cases}$
@@
Then $\tilde{\theta}>0$. Moreover, by the definition of $\tilde{\theta}$, for any $l\in\mathbb{N}$ the events $\{Z_{i}\leq 0\}$ and $\{Z_{i}\leq\tilde{\theta}/l\}$ are equivalent. Define the stopping time $T_{1}=\max\left\{j:T_{[0/n,j/n)}(W,0)\leq q_{\alpha}(j)\right\}.$ Set $I_{0}=[0/n,T_{1}/n)$. Then by definition of $T_{1}$ we have $T_{\mathring{I_{0}}}(W,0)\leq q_{\alpha}(T_{1})$. We defined the modification $\widetilde{W}_{i}=\mathbbm{1}_{\{Z_{i}\leq\tilde{\theta}\}}$ for all $i/n\in I_{0}$, then $W_{i}=\widetilde{W}_{i}$. Thus, $T_{\mathring{I_{0}}}(\widetilde{W},\tilde{\theta})=T_{\mathring{I_{0}}}(W,0)\leq q_{\alpha}(T_{1}),$ which indicates that $\tilde{\theta}$ fulfils the multiscale side-constrain on $\mathring{I_{0}}$ and therefore, we can estimate ${f}|_{I_{0}}$ with $\tilde{\theta}$. We continue this procedure until obtain an estimator for whole f{f}. More precisely, for any $k\in\mathbb{N}$, let
$T_{k}=\max\left\{j:T_{[T_{k-1}/n,j/n)}(W,0)\leq q_{\alpha}(j-T_{k-1})\right\},$
@@
and $I_{k-1}=[T_{k-1}/n,T_{k}/n)$. The restriction ${f}|_{I_{k-1}}$ can be estimated by $\tilde{\theta}/k$. Finally, let $\widetilde{K}=\max\{k:T_{k}<n\},$ and estimation is given by $\widetilde{f}=\sum_{k=0}^{\tilde{K}}\frac{\tilde{\theta}}{k+1}\mathbbm{1}_{I_{k}}.$ Since $\tilde{f}$ is piecewise different, it has exactly $\widetilde{K}$ change points. Thus, applying Proposition 7, we have $\mathbb{P}\left(\widehat{K}\geq k\right)\leq\mathbb{P}\left(\widetilde{K}\geq k\right)\leq\alpha^{k},$ and $\mathbb{E}\left(\widehat{K}\right)\leq\mathbb{E}\left(\widetilde{K}\right)\leq\frac{\alpha}{1-\alpha}.$
A.2 Proofs of Theorems 14
@@
$2(p-q)^{2}\leq p\log\frac{p}{q}+(1-p)\log\frac{1-p}{1-q}\leq\frac{(p-q)^{2}}{q(1-q)}.$
@@
The first inequality follows from Pinsker’s inequality (Tsybakov, , 2009, Lemma 2.5) and Vanegas et al., (2022, Theorem A.3). For the second one, by $x(x-1)\geq x\log(x)$ for $x>0$, we have
$\frac{p}{q}\log\frac{p}{q}\leq\frac{p}{q}\left(\frac{p}{q}-1\right)\quad\text{and}\quad\frac{1-p}{1-q}\log\frac{1-p}{1-q}\leq\frac{1-p}{1-q}\left(\frac{1-p}{1-q}-1\right).$
@@
$T_{I}(Z,\theta)\;=\;\underset{J\subseteq{I}}{\max}\sqrt{2L_{J}(Z,\theta)}-\sqrt{2\log\frac{e\lvert I\rvert}{\lvert J\rvert}}$,
with $L_{J}(Z,\theta)=\lvert J\rvert\bigl[\overline{W}_{J}\log(\frac{\overline{W}_{J}}{\beta})+(1-\overline{W}_{J})\log(\frac{1-\overline{W}_{J}}{1-\beta})\bigr],$ $\overline{W}_{J}=\left(\sum_{i\in J}W_{i}\right)/\lvert J\rvert$ and $W_{i}=\mathbbm{1}_{\{Z_{i}\leq\theta\}}\overset{i.i.d.}{\sim}{\rm Ber}(\beta)$ under $H_{I}$. Since $T_{I}(Z,\theta)$ is determined by $W$, in this proof we use $T_{I}(W)$ instead. For any $J\subseteq I$, define $a_{J}\in\mathbb{R}^{\lvert I\rvert}$ by
$a_{J,k}=\begin{cases}\frac{1}{\sqrt{\lvert J\rvert}},&\qquad\text{if }k\in J,\\ 0,&\qquad\text{otherwise},\end{cases}$
where $a_{J,k}$ denotes the $k$-th component of $a_{J}$. Set $A=\{a_{J}:J\subseteq I\}$ and $\lambda_{a}=\sqrt{2\log en\lVert a\rVert_{\infty}^{2}}$. We further define $R_{I}(W)$ and $S_{I}(W)$ by
$R_{I}(W)=\underset{a\in A}{\max}\sqrt{2}\lvert a^{\top}(W-\beta)\rvert-\lambda_{a}\quad\text{and}\quad S_{I}(W)=\underset{a\in A}{\max}\sqrt{\frac{2}{\beta(1-\beta)}}\lvert a^{\top}(W-\beta)\rvert-\lambda_{a}.$
Then, by Lemma 9, we have $R_{I}\leq T_{I}\leq S_{I}.$ For any $x,y\in\{0,1\}^{\lvert I\rvert}$ and weight $a\in\mathbbm{R}_{\geq 0}^{\lvert I\rvert}$, the weighted Hamming distance $d_{a}$ is given by $d_{a}(x,y)=\sum_{i=1}^{\lvert I\rvert}a_{i}\mathbbm{1}_{\{x_{i}\neq y_{i}\}}.$ We show that $T_{I}$ has the following Lipschitz property for the distance $d_{a}$ with Lipschitz constant $L=\sqrt{2}$: For any $x\in\{0,1\}^{\lvert I\rvert}$, there exists a $a=a(x)\in\mathbbm{R}_{\geq 0}^{\lvert I\rvert}$ with Euclidean norm $\lVert a\rVert=1$ such that $T_{I}(x)-T_{I}(y)\leq Ld_{a}(x,y),\text{ for all }y\in\{0,1\}^{\lvert I\rvert}$. For any $x,y\in\{0,1\}^{\lvert I\rvert}$, let $\tilde{a}=\tilde{a}(x)$ denote the maximizer of $S_{I}(x)$. Then
$T_{I}(x)-T_{I}(y)\leq S_{I}(x)-R_{I}(y)$
@@
where the second last equality follows since $x,y\in\{0,1\}^{\lvert I\rvert}$. Let $M_{I}$ denote the median of $T_{I}$. By Frick et al., (2014), Dümbgen and Spokoiny, (2001) and Dümbgen et al., (2006), the random sequence $T_{I}$ converges in distribution to an almost surely finite random variable $M$ as $\lvert I\rvert\to\infty$. Thus, median of $M$ is finite and $C\coloneqq\sup_{\lvert I\rvert\in\mathbb{N}}M_{I}<\infty$. Applying Giné and Nickl, (2016, Corollary 3.2.5), we have $\mathbb{P}(T_{I}-M_{I}\geq t)\leq 2\exp{\left(-t^{2}/8\right)}$ for any $t>0$. By choosing $t=2\sqrt{2\log(2/\alpha)}$, we have
$q_{\alpha}(\lvert I\rvert)\leq M_{I}+2\sqrt{2\log\frac{2}{\alpha}}\leq C+2\sqrt{2\log\frac{2}{\alpha}}$
for any $\lvert I\rvert\in\mathbb{N}$. ∎
@@
$I_{k}=\left(\tau_{k}-\tfrac{\lambda_{k}}{2},\tau_{k}+\tfrac{\lambda_{k}}{2}\right),$
@@
[…]
Assume that $\widehat{K}<K$ and let $\widehat{I}_{0},\dots,\widehat{I}_{\widehat{K}}$ denote the estimated segments. Since $\widehat{K}<K$, there exists a interval $I_{k}$ such that $I_{k}$ is completely contained in some estimated segment, say $\widehat{I}_{*}$. This can be verified by contradiction: Suppose that there is no such interval. Then any interval $I_{k}$ contains at least one estimated change point. Since $I_{k}$’s are disjoint, then we have $\widehat{K}\geq K$, which contradicts to the underestimation assumption. Thus, there exists a $I_{k}$ is contained in some estimated segment $\widehat{I}_{*}$, i.e., no change point is detected on $I_{k}$.
@@
[…]
@@
[…]
Obviously, we have $C_{k}\subseteq C_{k}^{+}\cup C_{k}^{-}$. Note that $C_{k}^{+}$ and $C_{k}^{-}$ are independent. We obtain $\mathbb{P}(C_{k})\leq 1-(1-\mathbb{P}(C_{k}^{+}))(1-\mathbb{P}(C_{k}^{-})).$ We provide an upper bound of $\mathbb{P}(C_{k}^{-})$, then the probability $\mathbb{P}(C_{k}^{+})$ can be upper bounded in a similar way. For any $i\in I_{k}^{-}$, we have $F_{i}^{-1}(\beta)=\theta_{k}^{-}$. Further set $F^{-}_{\min,k}=\min_{i\in I_{k}^{-}}F_{i}$ and $\bar{F}_{k}^{-}=\left(\sum_{i\in I_{k}^{-}}F_{i}\right)/\lvert I_{k}^{-}\rvert$. Let $\zeta_{I_{k}^{-}}=\inf\bigl\{x\in\mathbb{R}:\sum_{i\in I_{k}^{-}}\mathbbm{1}_{\{Z_{i}\leq x\}}\geq\lvert I_{s}^{-}\rvert\beta\bigr\}$ be the empirical quantile of the observations on $I_{k}^{-}$. For all $c\geq\theta_{k}^{-}+\delta_{k}/2$, if $\zeta_{I_{k}^{-}}\leq\theta_{k}^{-}+\delta_{k}/2$, then
@@
where $\overline{W}(Z,x)=\sum_{i\in I_{k}^{-}}\mathbbm{1}_{\{Z_{i}\leq x\}}/\lvert I_{k}^{-}\rvert$. The function
$g:x\mapsto\lvert I_{k}^{-}\rvert\left(x\log\frac{x}{\beta}+(1-x)\log\frac{1-x}{1-\beta}\right)$
is strictly convex with minimum $g(\beta)=0$, hence $g$ is strictly increasing on $(\beta,1)$. Thus, by (20) $L_{I_{k}^{-}}(Z,c)=g\left(\overline{W}(Z,c)\right)\geq g\left(\overline{W}\left(Z,\theta_{k}^{-}+\frac{\delta_{k}}{2}\right)\right)=L_{I_{k}^{-}}\left(Z,\theta_{k}^{-}+\frac{\delta_{k}}{2}\right).$ Therefore,
$\mathbb{P}(C_{k}^{-})\leq\mathbb{P}\bigl(C_{k}^{-}\cap\bigl\{\zeta_{I_{k}^{-}}\leq\theta_{k}^{-}+\frac{\delta_{k}}{2}\bigr\}\bigr)\\ +\mathbb{P}\bigl(\zeta_{I_{k}^{-}}>\theta_{k}^{-}+\frac{\delta_{k}}{2}\bigr).$ (21)
By Vanegas et al., (2022, Theorem A.2), the first term is bounded as
[…]
$\leq\mathbb{P}\left(L_{I_{k}^{-}}\left(Z,\theta_{k}^{-}+\frac{\delta_{k}}{2}\right)\leq\frac{1}{2}\left(q_{\alpha}(\lvert\widehat{I}_{*}\rvert)+\sqrt{2\log\frac{e\lvert\widehat{I}_{*}\rvert}{\lvert I_{k}^{-}\rvert}}\right)^{2}\right)$
$\leq 2\exp\left(-\frac{1}{2}\left(\sqrt{2n\lambda_{k}}\xi_{\bar{F}^{-}_{\beta}}(\frac{\delta_{k}}{2})-q_{\alpha}(\lvert\widehat{I}_{*}\rvert)-\sqrt{2\log\frac{e\lvert\widehat{I}_{*}\rvert}{\lvert I_{k}^{-}\rvert}}\right)^{2}_{+}\right).$
The second term in (21) is bounded as (via Vanegas et al., , 2022, Theorem A.3)
$\mathbb{P}\left(\zeta_{I_{k}^{-}}>\theta_{k}^{-}+\frac{\delta_{k}}{2}\right)\leq 2\exp\left(-n\lambda_{k}\xi_{F^{-}_{\min,\beta}}\left(\frac{\delta_{k}}{2}\right)^{2}\right).$
Note that the facts $\xi_{k}\leq\xi_{F^{-}_{\min,k},\beta}(\delta_{k}/2)\leq\xi_{\bar{F}_{k}^{-},{\beta}}(\delta_{k}/2)$, $q_{\alpha}(\lvert\widehat{I}_{*}\rvert)\leq\max_{m\leq n}q_{\alpha}(m)$ and $\lvert\widehat{I}_{*}\rvert\leq n$. Thus, we have $\mathbb{P}(C_{k})\leq 1-(1-\mathbb{P}(C_{k}^{+}))(1-\mathbb{P}(C_{k}^{-}))\leq 1-\gamma_{n,k}^{2},$ where
$\gamma_{n,k}=1-2\exp{\left(-n\lambda_{k}\xi_{k}^{2}\right)}-2\exp{\left(-\left(\sqrt{n\lambda_{k}}\xi_{k}-\frac{1}{\sqrt{2}}\max_{m\leq n}q_{\alpha}(m)-\sqrt{\log\frac{2e}{\lambda_{k}}}\right)_{+}^{2}\right)}.$
If all of these $K$ change-points are detected, then $\widehat{K}\geq K$,
@@
which proves the first claim. By the definitions of $\Lambda$ and $\Omega$ in (8), we have for all $k=1,\dots,K$,
$\gamma_{n,k}^{2}\geq\Bigg(1-2\exp\left(-\left(\sqrt{n}\Omega-\frac{1}{\sqrt{2}}\underset{m\leq n}{\max}q_{\alpha}(m)-\sqrt{\log\frac{2e}{\Lambda}}\right)_{+}^{2}\right)-2\exp\left(-n\Omega^{2}\right)\bigg)^{2}\eqqcolon\gamma_{n}.$
Thus, applying Bernoulli’s inequality, we have
$\mathbb{P}(\widehat{K}<K)$ $\leq 1-\gamma_{n}^{K}$
$\leq 4K\exp\left(-n\Omega^{2}\right)+4K\exp\left(-\left(\sqrt{n}\Omega-\frac{1}{\sqrt{2}}\underset{m\leq n}{\max}q_{\alpha}(m)-\sqrt{\log\frac{2e}{\Lambda}}\right)_{+}^{2}\right).\qed$